OpenStudy (countrygirl1431):

Solve the following equation. Then place the correct number in the box provided. 2(P + 1) > 7 + P

OpenStudy (anonymous):

2(P + 1) > 7 + P 2p+2 > 7+p -2 -2 2p= 5 + p -p -p 1p=5 p=5 @countrygirl1431

OpenStudy (countrygirl1431):

@paolaxx33 thank you so much for all your help

OpenStudy (anonymous):

No problem :) can you medal ? @countrygirl1431

OpenStudy (countrygirl1431):

already got ya

OpenStudy (countrygirl1431):

could you help me with a few more if u have time @paolaxx33

OpenStudy (anonymous):

Sure :) @countrygirl1431

OpenStudy (countrygirl1431):

Solve the following equation. Then place the correct number in the box provided. 2x < 6

OpenStudy (countrygirl1431):

i believ its 3

OpenStudy (anonymous):

you are correct @countrygirl1431

OpenStudy (countrygirl1431):

Solve the following inequality. Then place the correct answer in the box provided. Answer in terms of an mixed number. 8P + 2 > 3P - 15

OpenStudy (anonymous):

8P + 2 > 3P - 15 -2 -2 8p > 3p -17 -3p -3p 5p > -17 p=-3.4 @countrygirl1431

OpenStudy (countrygirl1431):

how would i write that in a fraction

OpenStudy (anonymous):

im not so sure

OpenStudy (countrygirl1431):

i have to fill in a box

OpenStudy (countrygirl1431):

would it help if i sent a screenshot? @paolaxx33

OpenStudy (anonymous):

sure send it

OpenStudy (countrygirl1431):

@paolaxx33